OpenStudy (anonymous):

which equation has a graph parallel to the graph of 9x + 3y = -22? y = 3x - 22 y = -3x + 8 y = 1/3(x) + 12 y = -(1/3)x - 2

OpenStudy (jack1):

parallel... so has same gradient y = mx + c gradient is m ur equation is 9x + 3y = -22 3y = -9x -22 y = -9/3x - 22/3 gradient = -9/3 = -3

OpenStudy (jack1):

which other answer has a gradient of -3?